OpenStudy (anonymous):

A (7, -3), B (4, y) the midpoint of AB is (1, 6), what is y?

OpenStudy (anonymous):

mid point formula is \[x =\frac{ x _{1} +x _{2} }{ 2 }\]

OpenStudy (anonymous):

and the same for y

OpenStudy (anonymous):

\[6=\frac{ -3+y }{ 2}\] \[12=-3+y\] \[12+3=y\] \[y=15\]

OpenStudy (anonymous):

Thank you so much for your help. I really appreciate it.

OpenStudy (anonymous):

Did you understand ?